OpenStudy (anonymous):

8x-4+3(x+7)=6x-3(x-3) for X

OpenStudy (anonymous):

Expand using distributive property. -4 + ( 3 x + 3 * 7 ) + 8x = (-3) (x-3) + 6x Remove unnecesssary parantheses. -4 + 3x +21 + 8x = -3x +9 +6x 11x + 17 = 3x + 9 8x = -8 x = -1 That's all ^_^